# Nonlinear Schrödinger Network

### v2

@@ 4.2 Learning Curve and Parameter Convergence
Figure 4: The convergence of $\alpha,\beta_{2},$ and $\gamma$ in the first four layers of the Nonlinear Schrödinger Network during training on the Starlight dataset. (a), (b), and (c) depict the convergence of $\alpha,\beta_{2},$ and $\gamma$, respectively. All parameters start with zero and gradually converge to values that optimize the network’s classification performance.
@@ 4.3 Ablation Study on the Impact of Dispersion and Nonlinearity
Baseline 49.1 (0.5) 15.6 (1.1)
$N$ Only 51.7 (1.2) 19.9 (1.9)
$D$ Only 79.5 (6.2) 60.3 (5.5)
Full 86.2 (3.3) 70.1 (2.5)
@@ 6 Conclusion
In this paper, we have presented the Nonlinear Schrödinger Network, a novel physics-inspired trainable model that leverages the principles of the Nonlinear Schrödinger Equation (NLSE) to learn complex patterns and nonlinear mappings from data. Unlike conventional neural networks, which often rely on large numbers of parameters, our approach directly utilizes the partial differential equation (PDE) that governs physical systems in a trainable and digital form, resulting in an interpretable and parameter-efficient model that can be optimized with backpropagation and gradient descent.
The experiments on various time series classification datasets have demonstrated that the Nonlinear Schrödinger Network achieves comparable or superior accuracy with orders of magnitude fewer parameters than conventional models such as Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). This highlights the efficiency of our approach in capturing complex dynamics with minimal model complexity. Furthermore, the transformations learned by the Nonlinear Schrödinger Network can be explicitly formulated and interpreted through the lens of a partial differential equation, with each parameter having a clear physical meaning. This contrasts sharply with the black-box nature of typical neural networks, offering insights into the data transformation process and making the model more transparent and understandable. While the current implementation uses the NLSE, the proposed method of using a physics equation as a trainable model to learn nonlinear mappings from data is not restricted to the NLSE but maybe implemented with other equations.

### v3

@@ 3.2 Learning Curve and Parameter Convergence
Figure 5: The convergence of $\alpha$ or attenuation, $\beta_{2}$ or group delay dispersion, and $\gamma$ or nonlinearity, of the physical system in the first four layers of the Nonlinear Schrödinger Network during training on the Starlight dataset. (a), (b), and (c) depict the convergence of $\alpha,\beta_{2},$ and $\gamma$, respectively. All parameters start with zero and gradually converge to values that optimize the network’s classification performance.
@@ 3.3 Ablation Study on the Impact of Dispersion and Nonlinearity
Baseline (Linear Classifier Only) 49.1 (0.5) 15.6 (1.1)
Nonlinearity ($N$) Only 51.7 (1.2) 19.9 (1.9)
Dispersion ($D$) Only 79.5 (6.2) 60.3 (5.5)
Full ($N$ and $D$) 86.2 (3.3) 70.1 (2.5)
@@
4 Extension to Gross-Pitaevskii Equation
The mathematical framework established for the Nonlinear Schrödinger Network can be generalized to other partial differential equations, demonstrating the broader applicability of the physical embedding paradigm. A natural extension is to the Gross-Pitaevskii equation (GPE), which serves as a mean-field approximation for the quantum dynamics of dilute boson gases, particularly in describing Bose-Einstein condensates at ultra-low temperatures [32]. The GPE is structurally very similar to the NLSE with the inclusion of an external potential term, thereby providing enhanced capabilities for modeling spatially confined quantum systems.
To explicitly establish the mapping between our NLSE implementation and the GPE framework, we formulate the Hamiltonian of the GPE as:
$i\hbar\frac{\partial\psi(\mathbf{r},t)}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{r})+g|\psi(\mathbf{r},t)|^{2}\right]\psi(\mathbf{r},t)$ (4)
When mapping between these equations, we establish the following parameter correspondences: the dispersion term $\beta_{2}$ in our NLSE framework corresponds to $-\hbar^{2}/m$ in the GPE formulation, while the nonlinearity coefficient $\gamma$ maps to $-g$, where $g$ characterizes interparticle interactions in the condensate. Conversely, the GPE introduces a potential term $V(\mathbf{r})$ that captures spatial confinement effects absent in the original NLSE. This term becomes an additional trainable parameter in our network architecture, complementing the existing parameter set and enhancing the model’s representational capacity. The implementation of the GPE within our network architecture requires a structural modification to the layer composition. Each Gross-Pitaevskii Layer is characterized by an augmented parameter set: $\{\beta_{2}^{(i)},\gamma^{(i)},V^{(i)}\}$, where $V^{(i)}$ parametrizes the external potential function. The potential transformation can be parameterized as:
$V^{(i)}(E(t))=E(t)\cdot\exp\left(-iV(t)\delta z\right)$ (5)
The potential function in this case is parameterized by a learned polynomial function, with polynomial order considered a hyperparameter, in our case kept at 20. Table 3 presents an ablation analysis comparing the performance metrics of various configurations of the GPE-based network against the baseline linear classifier. The results suggest that the inclusion of the potential term $V$ enhances the model’s capacity over the Nonlinear Schrödinger Network when the linear operation $D$ is included but the nonlinear $N$ is excluded. The dataset-specific performance improvements observed with the GPE implementation underscore the ability of the potential term to replace the nonlinearity term in the Nonlinear Schrödinger Network formulation.
Table 3: Ablation study on the impact of potential term from Gross-Pitaevskii equation.
Model Ford Engine Acc (%)
Baseline (Linear Classifier Only) 49.1 (0.5)
Nonlinearity ($N$) and Potential Energy ($V$) Only 52.1 (0.1)
Dispersion ($D$) and Potential Energy ($V$) Only 85.0 (7.6)
Full ($N$ and $D$ and $V$) 86.0 (2.4)
00footnotetext: Baseline: model with a linear classifier. Nonlinearity $N$ and Potential $V$ Only: GPE Network with only nonlinear and potential operations in each layer. Dispersion $D$ and Potential $V$ Only: GPE Network with only dispersive and potential operations in each layer. Full: Dispersive, nonlinear, and potential terms are all utilized.
@@ 5 Discussion
The experiments on various time series classification datasets have demonstrated that the Nonlinear Schrödinger Network achieves comparable or superior accuracy with orders of magnitude fewer parameters than conventional models such as Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). This highlights the efficiency of our approach in capturing complex dynamics with minimal model complexity. Furthermore, the transformations learned by the Nonlinear Schrödinger Network can be explicitly formulated and interpreted through the lens of a partial differential equation, with each parameter having a clear physical meaning. This contrasts sharply with the black-box nature of typical neural networks, offering insights into the data transformation process and making the model more transparent and understandable.
The extension to the GPE demonstrates that our approach is not limited to the NLSE but constitutes a general framework that can be adapted to various PDEs that model different physical phenomena. The success of this extension suggests promising avenues for incorporating other master equations of physics into trainable computational models.
5.1 Implementation in Analog Optical Computing
Given the physical underpinnings of the approach, there is potential for analog implementation, where the ultrafast timescales of pulse propagation can drastically reduce latency compared to traditional digital systems [33]. While constructing a system that implements a layer-by-layer NSN poses considerable engineering challenges, a new concept in AI optical hardware acceleration offers a promising alternative [20]. This method modulates data onto a supercontinuum laser spectrum, which then undergoes a nonlinear optical transformation analogous to a kernel operation, enhancing data classification accuracy. It is shown that the nonlinear optical kernel can improve linear classification results similar to a traditional numerical radial-basis-function (RBF) kernel but with orders of magnitude lower latency. The inference latency of the RBF kernel with a linear classifier is on the order of $10^{-2}$ to $10^{-3}$ second, while the nonlinear Schrödinger kernel achieves a substantially reduced latency on the order of $10^{-5}$ seconds and better classification accuracy.
In the original work, the performance is data-dependent due to the limited degrees of freedom and the unsupervised nature of the optical kernel. A follow-up work has demonstrated that by modulating the spectral phase of the input data, the nonlinear optical interactions within the kernel can be effectively trained in a data-driven manner [34]. This approach exploits a similar physical embedding process, with the same dispersive and nonlinear effects, reinterpreted as adjustable spectral-phase manipulations. Through this process, data-driven training and inference becomes feasible in a physically constrained, ultrafast domain. Such approaches provide potential groudwork for analog implementation of the Nonlinear Schrödinger Network.
5.2 Generalizability
@@
5.3 Other Limitations
A natural question is how broadly this idea applies to datasets or tasks that do not exhibit wave-like or dispersive phenomena. Since not all data domains align naturally with the physical assumptions underlying NLSE (or even PDEs in general), it raises questions about how effective “physical embedding” would be for, say, image-based tasks, high-dimensional tabular data, or domains with more complex spatiotemporal patterns. More varied benchmarking (particularly comparing to state-of-the-art methods on widely used benchmark corpora) would also help establish how well the approach generalizes beyond relatively moderate-sized data.
@@ 6 Conclusion
In this paper, we introduced the concept of physical embedding and demonstrated it through the Nonlinear Schrödinger Network (NSN), which repurposes the Nonlinear Schrödinger Equation (NLSE) as a trainable model for learning complex data mappings. Instead of relying on large sets of opaque parameters, the NSN embeds representations in a compact parameter space governed by physically interpretable coefficients. This approach, which directly encodes the data into a set of fundamental wave-equation properties, offers improved interpretability and drastically reduced memory requirements.
Our experiments on time series classification show that the NSN performs on par with, or surpasses, standard deep-learning architectures—yet requires orders of magnitude fewer parameters. Ablation studies highlight that its synergy of dispersion and nonlinearity is key to modeling intricate temporal patterns. Moreover, the concept of physical embedding readily adapts to other partial differential equations such as the Gross-Pitaevskii Equation, underscoring its broader relevance wherever domain-specific physics can be leveraged.
In addition to these gains in parameter efficiency, analog implementation stands out as an exciting future direction. By physically realizing dispersion and nonlinear effects in hardware—particularly in optical platforms—the NSN could perform ultrafast inference with significantly lower energy consumption. While there are trade-offs in terms of domain specificity and flexibility, the prospect of physical embedding offers a powerful alternative to black-box networks, promising deeper interpretability, resource-conscious scaling, and new pathways for physics-based machine learning.